OpenStudy (anonymous):

Simplify: 9x^5 + 3x^2 + 6x over 3x

OpenStudy (whpalmer4):

\[\frac{a+b+c}{d} = \frac{a}{d} + \frac{b}{d} + \frac{c}{d}\] You can use that to rewrite your problem from \[\frac{9x^5+3x^2+6x}{3x}\]to\[\frac{9x^5}{3x} + \frac{3x^2}{3x} + \frac{6x}{3x}\] Do you know how to simplify each of those terms?

OpenStudy (anonymous):

I don't.

OpenStudy (whpalmer4):

Can you at least reduce the numeric part of the fractions?

OpenStudy (whpalmer4):

If I asked you to write the fractions \[\frac{9}{3}, \,\frac{3}{3},\, \frac{6}{3}\]in their simplest form, what would your answer be?

OpenStudy (anonymous):

No.. I'm really quite pathetic when it comes to math.

OpenStudy (whpalmer4):

You can't divide 9 by 3? 3 by 3? 6 by 3?